# Towards Real World Debiasing: A Fine-grained Analysis On Spurious Correlation

### v2

@@ Appendix F Related Works
Generally, spurious correlation refers to the phenomenon that two distinct concepts are statistically correlated within the training distribution, though there is no causal relationship between them, e.g. background and foreground object [49]. The spurious correlation is a vital aspect of understanding how machine learning models learn and generalize [43]. Specifically, studies on distribution shift [6] claim that spurious correlation is one of the major types of distribution shift in the real world, and thus an important distribution shift that a reliable model should be robust to. Furthermore, studies on fairness and bias [50] have demonstrated the pernicious impact of spurious correlation in classification [34], conversation [51], and image captioning [22]. However, despite its broad impact, spurious correlation is generally used as a vague concept in previous works and lacks a proper definition and deeper understanding of it. This is also the major motivation of this work.
Spurious correlation debiasing.
In this work, we focus only on debias methods that do not require bias information, i.e. without annotation on the spurious attribute, for it is more practical. The mainstream of work [12, 9, 15, 52, 17, 18, 20, 53] in the area of spurious correlation debiasing follows the framework proposed by Nam et al. [12], which introduces a biased auxiliary model to capture biases within the training data, according to which a re-weighing scheme designed to emphasize BC samples is applied to train a debiased model, i.e. DBAM methods. Based on the framework, Lee et al. [9] further proposed a feature augmentation technique to further enhance BC samples. Hwang et al. [52] proposed to augment biased data by applying mixup [54] to contradicting pairs. Lim et al. [17] proposed an adversarial-based approach to augment BC samples aiming to increase the diversity of BC samples. Lee et al. [20] proposed to first filter out BC samples before training the biased model aiming to enhance the bias capture process of the biased model. Liu et al. [16] regard the samples misclassified by the model trained with empirical risk minimization as BC samples and emphasize them during training of a debiased model. Recently, Park et al. [53] proposed to provide models with explicit spatial guidance that indicates the region of intrinsic features.
Among other lines of work, Park et al. [53] create images without bias attributes using an image-to-image translation model [55]. A recent pair-wise debiasing method $\chi^{2}$ model [9] encourages the model to retain intra-class compactness using samples generated via feature-level interpolation between BC and BA samples.
@@ Appendix G Limitations and Future Work
We uncover the insufficiency of existing debiasing benchmarks theoretically and empirically, highlighting the importance of debiasing on real-world biases. We further proposed a feature-destruction-based method that focuses on DBAM methods. However, there are still a few limitations of this work:
  • While DBAM methods are the predominant works in debiasing, there are also other lines of work such as data generation methods. Thus, one limitation is that We have not evaluated such methods with our proposed evaluation framework which might also yield some interesting insights on debiasing.
  • Another limitation is that, though we have already seen the potential of target feature destruction, whether it can be applied to other lines of work remains to be studied.
  • As shown in section E, while our proposed approach effectively improves the performance of existing DBAM methods on all biased distributions from the real world, the performance is still far from satisfactory, which remains to be further improved in future works.
We see potential within those limitations and leave them for future research.
@@ Appendix H Boarder Impact
From a technical standpoint, our research provides a comprehensive framework for analyzing and mitigating biases in datasets. The proposed fine-grained analysis framework and evaluation benchmarks offer a new perspective on how biases manifest in real-world data and how existing debias methods can be improved. Our approach, which involves the destruction of target features during bias capture, demonstrates significant improvements in handling real-world biases, as evidenced by our extensive experimental results.
By advancing the understanding of dataset biases and improving the performance of debias methods, our research contributes to the development of more robust and generalizable AI models. This is particularly relevant in an era where AI systems are increasingly deployed in dynamic and diverse environments, necessitating models that can adapt and maintain high performance across different contexts and populations.

### v3

@@ Appendix F Related Works
Generally, spurious correlation refers to the phenomenon that two distinct concepts are statistically correlated within the training distribution, though there is no causal relationship between them, e.g. background and foreground object [chu2024task]. The spurious correlation is a vital aspect of understanding how machine learning models learn and generalize [48]. Specifically, studies on distribution shift [6] claim that spurious correlation is one of the major types of distribution shift in the real world, and thus an important distribution shift that a reliable model should be robust to. Furthermore, studies on fairness and bias [53] have demonstrated the pernicious impact of spurious correlation in classification [36], conversation [54], and image captioning [23]. However, despite its broad impact, spurious correlation is generally used as a vague concept in previous works and lacks a proper definition and deeper understanding of it. This is also the major motivation of this work.
Debiasing without bias supervision.
In this work, we focus only on debiasing methods that do not require bias information, i.e. without annotation on the spurious attribute, for it is more practical. Existing work [12, 9, 15, 55, 17, 18, 20, 56] in the area generally involve a biased auxiliary model to capture biases within the training data, according to which the debiased is trained with various techniques. We call such paradigm debiasing with biased auxiliary model (DBAM). Specifically, Nam et al. [12] is the first work that follows the DBAM paradigm, proposing to use GCE for bias capture, and the loss-based sample re-weighing scheme to train the debiased model. Lee et al. [9] further proposed a feature augmentation technique to further utilize the captured bias, enhancing the BC samples. Hwang et al. [55] proposed to augment biased data identified according to the biased auxiliary model by applying mixup [57] to contradicting pairs. Lim et al. [17] proposed to conduct adversarial attacks on the biased auxiliary model to augment BC samples aiming to increase the diversity of BC samples. Lee et al. [20] proposed to first filter out BC samples before training the biased auxiliary model aiming to enhance the bias capture process of the biased model. Liu et al. [16] regard the samples misclassified by the biased auxiliary model as BC samples and emphasize them during training of the debiased model. Recently, Park et al. [56] proposed to provide models with explicit spatial guidance that indicates the region of intrinsic features according to a biased auxiliary model. Kim et al. [14] create images without bias attributes using an image-to-image translation model [58] built upon a biased auxiliary model. A recent pair-wise debiasing method $\chi^{2}$ model [59] based on biased auxiliary models encourages the debiased model to retain intra-class compactness using samples generated via feature-level interpolation between BC and BA samples.
@@ Appendix G Limitations and Future Work
We uncover the insufficiency of existing debiasing benchmarks theoretically and empirically, highlighting the importance of debiasing on real-world biases. We further proposed a feature-destruction-based method that focuses on DBAM methods. However, there are still a few limitations of this work: As shown in section E, while our proposed approach effectively improves the performance of existing DBAM methods on all biased distributions from the real world, the performance is still far from satisfactory, which remains to be further improved in future works. We see potential within those limitations and leave them for future research.
@@ Appendix H Boarder Impact
From a technical standpoint, our research provides a comprehensive framework for analyzing and mitigating biases in datasets. The proposed fine-grained analysis framework and evaluation benchmarks offer a new perspective on how biases manifest in real-world data and how existing debiasing methods can be improved. Our approach, which involves the destruction of target features during bias capture, demonstrates significant improvements in handling real-world biases, as evidenced by our extensive experimental results.
By advancing the understanding of dataset biases and improving the performance of debiasing methods, our research contributes to the development of more robust and generalizable AI models. This is particularly relevant in an era where AI systems are increasingly deployed in dynamic and diverse environments, necessitating models that can adapt and maintain high performance across different contexts and populations.